### v3

# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse Multi-Baseline Data

Abstract
Multi-baseline Synthetic Aperture Radar (SAR) three-dimensional (3D) tomography is a crucial remote sensing technique that provides 3D resolution unavailable in conventional SAR imaging. However, achieving high-quality imaging typically requires multi-angle or full-aperture data, resulting in significant imaging costs. Recent advancements in sparse 3D SAR, which rely on data from limited apertures, have gained attention as a cost-effective alternative. Notably, deep learning techniques have markedly enhanced the imaging quality of sparse 3D SAR. Despite these advancements, existing methods primarily depend on high-resolution radar images for supervising the training of deep neural networks (DNNs). This exclusive dependence on single-modal data prevents the introduction of complementary information from other data sources, limiting further improvements in imaging performance. In this paper, we introduce a Cross-Modal 3D-SAR Reconstruction Network (CMAR-Net) to enhance 3D SAR imaging by integrating heterogeneous information. Leveraging cross-modal supervision from 2D optical images and error transfer guaranteed by differentiable rendering, CMAR-Net achieves efficient training and reconstructs highly sparse multi-baseline SAR data into visually structured and accurate 3D images, particularly for vehicle targets. Extensive experiments on simulated and real-world datasets demonstrate that CMAR-Net significantly outperforms state-of-the-art sparse reconstruction algorithms based on compressed sensing (CS) and deep learning (DL), with average improvements of 75.83% in PSNR and 47.85% in SSIM. Furthermore, our method eliminates the need for time-consuming full-aperture data preprocessing and relies solely on computer-rendered optical images, significantly reducing dataset construction costs. This work highlights the potential of deep learning for multi-baseline SAR 3D imaging and introduces a novel framework for radar imaging research through cross-modal learning.
keywords:
Multi-baseline SAR , Sparse imaging , 3D reconstruction , Cross-modal learning , Information fusion
[…]
[bit-phy2]organization=Beijing Key Laboratory of Nanophotonics and Ultrafine Optoelectronic Systems, School of Physics, Beijing Institute of Technology,addressline=No.5 Zhongguancun South Street, city=Haidian District, postcode=100081, state=Beijing, country=China
1 Introduction
Multi-baseline SAR 3D tomography is a highly promising technique for SAR-based 3D reconstruction, due to its ability to provide complete threedimensional spatial resolution[1]. Multi-baseline SAR forms a synthetic aperture in the elevation dimension by performing multiple flights at varying altitudes, providing height resolution, enabling 3D imaging. This technology has significant potential in applications such as high-precision geographic remote sensing, urban 3D mapping, and detailed target interpretation [2]. In contrast, traditional 2D SAR generates a two-dimensional projection of a 3D target in the slant-range azimuth plane, leading to distortions such as overlap and perspective shortening, which result in poor image clarity and limited interpretability [3]. SAR 3D imaging, on the other hand, captures the spatial distribution of scattering centers in three dimensions, providing an accurate and intuitive representation of the target’s structure. As a result, it has gradually become a key focus of SAR technology research.
However, applying multi-baseline SAR for 3D observation presents several challenges in practical scenarios. To obtain a 3D image, multi-baseline SAR requires observations from multiple long baselines or even circular trajectories around the target area to achieve high angular resolution[4, 5, 6]. Due to constraints such as terrain, weather, and flight conditions, observations are often limited to a few angles or apertures along restricted flight paths, making it difficult to meet the data requirements. Moreover, the large volume of data generated by multi-baseline observations places significant demands on data computation and storage [7]. To address these issues, researchers have focused on developing sparse imaging algorithms that can reconstruct detailed 3D target images from limited and incomplete data. Existing methods can generally be classified into two main categories: those based on compressed sensing and those based on deep learning.
Compressed sensing (CS) has long been the mainstream method for reconstructing signals from incomplete and sparse measurements, and it was applied early in multi-baseline SAR 3D imaging [8, 9]. These methods treat the sparse multi-baseline SAR 3D imaging problem as a signal recovery model. By introducing various penalty functions and optimization techniques, they are able to reconstruct high-resolution target images from incomplete measurements. Typical works, such as [9, 10], have reduced the reliance of multi-baseline SAR 3D imaging on complete measurements. However, CS-based methods are computationally intensive, especially when dealing with 3D data. Additionally, CS solvers are highly sensitive to parameter settings, which limits their generalization and hampers their further development and application[11, 12].
Deep learning (DL) techniques have been increasingly applied in the field of sparse multi-baseline SAR 3D imaging in recent years. Researchers have explored deep neural networks (DNNs) to learn image priors from training data and apply them for sparse imaging[13, 14, 15, 16, 17]. Thanks to their parallel structure and inference-based (non-iterative) mode, DL methods offer more efficient and stable imaging results compared to CS-based approaches[18]. Based on their implementation, DL-based methods can be divided into two types.
The first type of method embeds NN models into the CS framework, replacing the traditional iterative process with cascaded DNNs that model the nonlinear optimization components. Notable works include [13, 19, 20, 21]. After extensive training, the network learns data priors and optimization parameters, eliminating the need for manual configuration. Unlike CS-based methods, this approach requires only a single inference to achieve reconstruction accuracy comparable to CS algorithms, significantly reducing the computational complexity.
The second type of method combines traditional imaging algorithms with DNNs. It first applies traditional algorithms to pre-imaging sparse data, and then uses DNNs to refine the results. For example, in [22], a 3D U-Net was used to improve back-projection (BP) pre-imaging, enabling high-resolution reconstruction of distorted BP images. Similarly, in [23], a GAN-based sparse-aspects-completion network (SACNet) enhanced the results of CS pre-imaging, demonstrating that a network trained on simulated data can still achieve good target reconstruction on real-world data. These methods frame the improvement of pre-imaging results as an image enhancement or denoising task. Thanks to the strong data representation ability of DNNs and extensive training, they can quickly and stably reconstruct high-resolution 3D target images with a single network inference, even under conditions of highly sparse observation data. As a result, these methods have become the state-of-the-art (SOTA) for sparse SAR 3D reconstruction.
In summary, DL-based sparse multi-baseline SAR 3D reconstruction methods demonstrate significant potential in enhancing both imaging efficiency and quality[18]. However, current learning-based approaches still face several significant practical challenges.
  1. 1.
    DL-based methods learn the mapping between training data pairs, which are derived from a single modality (electromagnetic information). This constraint on resolution, imposed by the electromagnetic nature, limits further improvements in image quality.
  2. 2.
    As a data-driven technology, DL-based methods rely heavily on the quality of the training data. Since high-resolution 3D SAR images are required for supervision, existing methods often necessitate time-consuming preprocessing of full-aperture data. This process consumes significant storage space and demands high-performance real-time processing platforms, making it costly and limiting the practical deployment of such algorithms.
  3. 3.
    The pre-imaging results fed into the network for high-resolution reconstruction are sensitive to the observation geometry of multi-baseline SAR, including the varying scattering characteristics of anisotropic targets at different angles, and are highly susceptible to noise interference. As a result, the reconstruction network must have strong generalization capabilities.
To overcome the aforementioned limitations, particularly the resolution constraints caused by single-modality information, in this paper, we introduce cross-modal learning. The cross-modal learning explores the intrinsic relationships between different data modalities, seeks shared representations, and integrates heterogeneous information. This concept has recently been successfully applied to problems such as SAR image translation[24, 25, 26].
Here, we propose a sparse multi-baseline SAR 3D reconstruction method based on cross-modal supervision and design a CMAR-Net to significantly enhance the quality of 3D SAR images. The supervisory information for training CMAR-Net is directly derived from aligned optical 2D images. The primary challenge of using 2D images to supervise a 3D imaging network lies in the difference in data formats. To address this, we employ differentiable rendering techniques to compute the consistency error between 2D pixels and 3D voxels, ensuring that this error can be back-propagated through the network. The main contributions of this paper are as follows:
  1. 1.
    To the best of our knowledge, this work is the first to introduce the concept of cross-modal learning into SAR 3D reconstruction. This approach enables high-resolution reconstructions with complete visual structure and distinct features, overcoming the resolution constraints inherent to electromagnetic images.
  2. 2.
    We propose a novel CMAR-Net that integrates a unique data augmentation strategy and a projection-reprojection module, enhancing both robustness and generalization. The network is trained on a limited set of simulated data and successfully validated on real-world data without any fine-tuning.
  3. 3.
    The proposed method only requires 2D optical images as supervisory data, eliminating the need for high-resolution images derived from full-aperture data. This simplifies the construction of high-quality imaging datasets, facilitating the broader application of 3D SAR reconstruction technology.
  4. 4.
    Extensive experiments demonstrate that our method achieves outstanding 3D target reconstruction results even under low signal-to-noise ratio (SNR) conditions and with highly sparse angular measurements. Image quality improves 75.83% in PSNR and 47.85% in SSIM compare to SOTA methods.
2 Methodology
2.1 Method framework
The proposed multi-baseline SAR 3D reconstruction method using sparse data is shown in Figure 1. Our method includes two modules: the pre-imaging module and the cross-modal reconstruction module. The pre-imaging module processes sub-aperture data and combines them noncoherently to produce a coarse, incomplete 3D image. The cross-modal reconstruction module, which includes a DNN trained with optical image supervision, processes the pre-imaging result to generate realistic 3D image with clear contours and complete structures. The details of each module are explained below.
Figure 1: The framework of the proposed sparse multi-baseline SAR cross-modal reconstruction method.
2.2 Pre-imaging module
Figure 1(left) illustrates the multi-baseline SAR multi-angle observation geometry. The radar platform performs multi-angle measurements around the target, shown as colored line segments. At each sub-angle, multiple straight-line flights at different altitudes generate synthetic apertures in both the azimuth and elevation dimensions, represented by dot clusters. Under far-field conditions, frequency-domain echo data are collected at different azimuth angles and heights for each sub-aperture, forming a 3D circular k-space observation of the target. After interpolation and coordinate transformation, the spatial positions of the target’s scattering centers can be obtained through a 3D inverse Fourier transform[8].
In the pre-imaging stage, phase errors from trajectory inaccuracies are corrected first, and the frequency-domain data in polar coordinates are then interpolated into Euler coordinates. A 3D inverse Fourier transform then reconstructs the target’s 3D image for each sub-aperture. Finally, the images from all sub-apertures are non-coherently summed to generate the pre-imaging result, which serves as input to the cross-modal reconstruction module.
2.3 Cross-modal reconstruction module
The cross-modal reconstruction module takes pre-imaging result as input and uses CMAR-Net to generate high-resolution 3D images that accurately represent the target’s physical shape. As shown in Figure 2, the network follows an encoder-decoder architecture with skip connections. The encoder extracts structural features from the incomplete pre-imaging result, while the decoder reconstructs the vehicle’s 3D image from the low-dimensional latent representation. To further improve feature representation, a projection-reprojection (PRP) unit is added near the bottleneck layer. At the output, a differentiable rendering module is introduced to enable cross-modal supervision. The following sections describe each module in detail.
2.3.1 Network architecture
Figure 2 illustrates the detailed structure and training process of the CMAR-Net. The network consists of a contraction path (on the left), an expansion path (on the right), and a PRP unit. The contraction path includes four downsampling layers, each consisting of a 3D convolutional layer followed by a max-pooling layer. The convolutional layers apply LeakyReLU activation and progressively increase the number of feature channels, while the pooling layers reduce the data dimensionality. The contraction path ends with a connection to the PRP unit.
The expansion path mirrors the contraction path, consisting of four upsampling layers and a single convolutional layer. Each upsampling layer uses a 3D transposed convolution with ReLU activation. Feature maps from the expansion path are combined with those from the contraction path at the same resolution through skip connections before being passed through the upsampling layers. These layers reduce the number of feature channels while increasing spatial resolution. The final upsampling layer produces a 64-channel output, then refined by a output convolution layer to smooth and integrate information across channels, ensuring a more refined reconstruction.
Figure 2: Network Architecture.
2.3.2 Projection-Reprojection unit
We design a PRP unit between the encoder and decoder to improve representation and generalization.
The structure of the PRP unit is detailed in Figure 3. The projection part consists of two fully connected layers. The first fully connected layer, activated by LeakyReLU, compresses the input feature map into a low-dimensional vector. The second fully connected layer eliminates the activation and further compresses the feature’s dimensionality to produce the latent representation $z$. The reprojection part reverses the projection process by symmetrically expanding the feature vector and reshaping it back to the original feature map size. The only difference is that the final fully connected layer uses ReLU activation to ensure consistency in the data dynamics within the network.
Figure 3: Architecture of PRP unit.
2.4 Differentiable volume rendering
Optical images contain rich, intuitive structural information, and, due to their passive imaging mechanism, high-resolution optical images are often easier to obtain in practical applications compared to radar electromagnetic images. However, integrating the advantages of 2D optical images into 3D SAR imaging requires bridging the dimensional gap between the two data types.
Differentiable volume rendering offers a promising solution. This technique converts 3D volumetric data into 2D images, and its differentiable nature allows gradient-based optimization of 3D structures using a 2D loss function, facilitating effective supervision of 3D reconstruction through 2D images. In recent years, it has been widely used in computational graphics[27], novel view synthesis[28] and point cloud completion[29, 30]. In our network, we incorporate a differentiable volume rendering module at the end of the CMAR-Net to generate 2D images from multiple views. By supervising the network with optical images, we leverage cross-modal benefits to improve the quality of the 3D SAR reconstruction.
Figure 4 illustrates the differentiable volume rendering process. Given a 3D volumetric data $\mathbf{V}\in\mathbb{R}^{W\times H\times D}$, a camera position $o$, and a viewing direction $\mathbf{d}$, volume rendering computes the pixel values $C(\mathbf{r})$ of a camera ray $\mathbf{r}(t)=o+t\mathbf{d}$ with near and far bounds $t_{n}$ and $t_{f}$ using the following formula:
[…]
Here, $\sigma(\mathbf{r}(t))$ denotes the volume density along the camera ray at point $\mathbf{r}(t)$ in the viewing direction $\mathbf{d}$, and $dt$ represents the step size of the ray at each integration step. $T(t)$ is the cumulative transmittance, which indicates the probability that the ray propagates from $t_{n}$ to $t$ without being intercepted. The $T(t)$ is given by the following equation:
[…]
Figure 4: Rendering diagram.
To compute the continuous integral, we discretize the ray by sampling it at equidistant depths and apply the integration rules discussed in the literature [31] to estimate the pixel value $C(r)$. This is given by the following formula:
$\hat{C}(\mathbf{r})=\sum\limits_{i=1}^{N}T_{i}(1-\text{exp}(-\sigma_{i}\delta_{i})).$ (3)
Where
$T_{i}=\text{exp}\left(-\sum\limits_{j=1}^{i-1}\sigma_{j}\delta_{i}\right).$ (4)
The $\delta_{i}=t_{i+1}-t_{i}$ represents the distance between adjacent samples. The final rendered image is obtained by concatenating the estimated pixel values along all camera rays passing through the pixels of the rendering canvas, as expressed by:
$I=\left[\begin{array}[]{ccc}\hat{C}(\mathbf{r}_{0,0})&\dots&\hat{C}(\mathbf{r}_{w-1,0})\\ \vdots&\ddots&\vdots\\ \hat{C}(\mathbf{r}_{0,h-1})&\dots&\hat{C}(\mathbf{r}_{w-1,h-1})\end{array}\right].$ (5)
$\mathbf{r}_{i,j}$ represents the camera ray passing through the pixel at $(i,j)$, where $i\in[0,h-1]$ and $j\in[0,w-1]$, with $h$ and $w$ denoting the height and width of the rendered image, respectively.
2.5 Loss design
We use images rendered from the target’s 3D model at different viewpoints to serve as ground truth to compute the reconstruction error. In this paper, we employ the Huber loss function to quantify the error. The Huber loss between the $i$-th viewpoint’s reconstructed image, $I_{r}^{i}$, and the corresponding ground truth image, $I_{g}^{i}$, is defined as:
$\mathcal{L}_{{\rm{huber}}}^{i}({I_{g}^{i}},{I_{r}^{i}},\gamma)=\left\{\begin{array}[]{l}{\textstyle{1\over 2}}{({I_{g}^{i}}-{I_{r}^{i}})^{2}},{\rm{\quad\quad\quad\thinspace\thinspace}}\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|\leq\gamma\\ \gamma\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|-{\textstyle{1\over 2}}{\gamma^{2}}.{\rm{\quad}}\left|{{I_{g}^{i}}-{I_{r}^{i}}}\right|>\gamma\end{array}\right.$ (6)
Here, $\gamma$ is a adjustable parameter of the Huber loss. For $V$ rendering viewpoints, the total loss function is defined as their average:
$\mathcal{L}=\frac{1}{V}\sum\limits_{i=0}^{V-1}{\mathcal{L}_{{\rm{huber}}}^{i}({I_{g}^{i}},{I_{r}^{i}},\gamma)}.$ (7)
3 Experiments
3.1 Experimental Settings
To train CMAR-Net, we construct a civilian vehicle dataset, including simulated sparse multi-baseline SAR 3D pre-images and multi-view optical images. For validation, we perform extensive comparative experiments with baseline methods on the simulated data, including feature space interpolation. We then validate the generalization performance of CMAR-Net using real-world data and conduct ablation studies on the PRP unit and data augmentation strategy. This section provides a detailed overview of the experiment setup.
3.1.1 Dataset and Augmentation
The sparse multi-baseline SAR 3D pre-images in the dataset were generated using the Civilian Vehicle Radar Data Domes (CVDomes), published by the United States Air Force Laboratory[32]. This data includes full-polarization, far-field X-band electromagnetic scattering data for ten civilian vehicles, covering a 360° azimuth and a pitch angle range from 30° to 60°. The viewpoint distribution and parameter settings of the simulation scene are illustrated in Figure 5 and Table 1. For the training set, data from five vehicles—two sedans, two SUVs, and one pickup truck—are selected. Imaging is performed at eight pitch angles between 44.25° and 46°, with a spacing of 0.1875° between each angle. The 360° azimuth data are divided into 72 sub-apertures, each 5° wide and spaced 0.0625° apart. Sub-aperture imaging is then performed, and finally, 8 randomly selected sub-apertures are non-coherently summed to generate the 3D pre-images for training.
Table 1: Simulated Parameters of the Civilian Vehicle Radar Data Domes[32].
Parameter Value
Radar center frequency 9.6GHz
Unambiguous range $\approx 15m$
Extrapolation extent $\leqslant 0.25^{\circ}$
Azimuth extent $360^{\circ}$
Elevation extent $30^{\circ}-60^{\circ}$
The multi-view optical images used for supervision in the dataset are generated by rendering 3D digital models of the vehicles. We collect 3D models of the same vehicle types as those in the CVDomes and rendered optical binary images from multiple viewpoints. The process of assembling the dataset is illustrated in Figure 5.
Additionally, we design a data augmentation strategy to expand our dataset. As shown in Figure 6, the augmentation involves three operations: scaling $S$, rotation $R$, and translation $T$. For each pair of input raw 3D pre-images, a geometric transformation is applied, which is also performed on the corresponding 3D digital model. This transformation is then reflected in the rendered images. This strategy increases the sampling density of the data space, which enhances the network’s generalization.
Figure 5: Dataset construction.
Figure 6: Data augmentation strategy.
3.1.2 Implementation
In the pre-imaging stage, the imaging scene is set with dimensions of $3.2m\times 3.2m\times 6.4m$ and a spatial resolution of $0.05m$, resulting in a data size of $64\times 64\times 128$. The pre-imaging data is then normalized to a dynamic range of [0, 1] before being input into the network. For the differentiable volume rendering module, the rendered image size is set to 256 $\times$ 256, with the camera positioned 7m from the scene center. The camera is fixed to point toward the center from 8 predetermined angles.
During training, the hyperparameters are set as follows: Huber loss parameter $\gamma=0.7$, batch size = 1, and the Adam optimizer with momentum parameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$. The initial learning rate is set to $1\times 10^{-4}$, gradually decreasing to $5\times 10^{-5}$. The implementation is based on the PyTorch framework, with training and inference conducted on a machine equipped with an NVIDIA RTX A6000. The network is trained and evaluated on the simulated data, and, without any further fine-tuning, it is ultimately validated on the real-world data.
3.1.3 Baseline methods
To demonstrate the improvement in imaging quality achieved by integrating optical information, we use traditional imaging methods and DL methods using radar image supervision as baseline comparisons. The traditional methods include BP[33] and CS[34], while the DL methods include SACNet[23] and UNet3D[22], which are specifically designed to enhance the quality of sparse multi-baseline SAR 3D imaging. Moreover, we use the backbone networks of SACNet and UNet3D, augmented with cross-modal (CM) supervision, as two additional baseline methods.
3.1.4 Evaluation Metrics
We convert the vehicle’s 3D digital model into volume data with the same spatial extent and resolution as the imaging settings, using it as the ground truth. The imaging quality of our method is then evaluated using full-reference metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Intersection over Union (IoU), and Cross-Entropy (CE).
3.2 Results for Simulated Data
3.2.1 Pre-imaging and Reconstruction Results
First, we provide an overview of the 3D reconstruction performance of CMAR-Net. Figure 7 illustrates the pre-imaging results fed into the network and the reconstruction results produced by the CMAR-Net. As shown in the figure, CMAR-Net significantly enhances the reconstruction of the vehicle’s surface, resulting in a reconstructed image with visually distinguishable vehicle features that closely resemble the ground truth (GT).
Figure 7: Overview of 3D reconstruction performance of CMAR-Net. Pre-imaging results (first row), CMAR-Net reconstructions (second row) and ground truth (third row). Imaging at an SNR of 30dB and 8 sub-apertures.
3.2.2 Comparison with Baseline Methods
Next, we compare CMAR-Net with baseline methods. Figure 8 presents the imaging results of different methods at an SNR of 30dB and 8 sub-apertures. Compared to the pre-imaging results in Figure 7, the BP method preserves the strong scattering points of the vehicle body by non-coherently integrating the maximum amplitude values from sub-aperture images. For instance, the bucket structure of Pickup are prominently improved. However, the imaging quality is degraded by high side lobes, as evidenced by the distorted front and windshield of the MPV and Camry. The CS method mitigates the impact of side lobes by optimizing sub-aperture imaging. However, the imaging results reveal significant structural omissions, attributed to the sparsity assumptions and the configuration of the optimization parameters.
DL methods supervised by radar images leverage prior knowledge from full-aperture imaging, enabling the reconstruction of vehicle images with more scattering details using sparse data. Such as the clearer bucket on the Pickup. However, due to constraints imposed by electromagnetic characteristics, the imaging results lack well-defined planar structures on the vehicle body.
In contrast, CMAR-Net does not suffer from structural loss and provides accurate geometry and clear vehicle contours, making it the best-performing method among all approaches. For example, in the reconstruction of the MPV, CMAR-Net restores a more complete rear body, while for the Jeep, it presents clearer contours and more precise geometric representation. Additionally, we found that after integrating cross-modal supervision, the performance of SACNet and UNet3D also improved significantly, further validating the applicability of the cross-modal supervision strategy to general networks.
Figure 8: Imaging results of different methods. Each column corresponds to an imaging method, and each row represents a vehicle type. Imaging at an SNR of 30dB and 8 sub-apertures.
To evaluate the reconstruction performance of CMAR-Net under different conditions, we conduct experiments where the SNR decreased from 25 dB to 5 dB, and the number of sub-apertures reduced from 12 to 4. Figures 9 to 13 present the imaging results for all experimental conditions.
As the SNR decreases and the number of sub-apertures reduces, the imaging results of the BP and CS methods show more side lobes, leading to reduced vehicle recognizability. SACNet and UNet3D demonstrate strong robustness under deteriorating imaging conditions, providing consistent reconstruction results. After integrating the cross-modal supervision, their performance remains stable. In contrast, CMAR-Net consistently delivers the best imaging performance across all vehicle types. Even under the most challenging conditions (5 dB SNR and 4 sub-apertures), CMAR-Net achieves excellent reconstruction, while the performance of other methods significantly deteriorates.
Figure 9: Imaging results of different methods at an SNR of 25dB and 12 sub-apertures.
Figure 10: Imaging results of different methods at an SNR of 20dB and 10 sub-apertures.
Figure 11: Imaging results of different methods at an SNR of 15dB and 8 sub-apertures.
Figure 12: Imaging results of different methods at an SNR of 10dB and 6 sub-apertures.
Figure 13: Imaging results of different methods at an SNR of 5dB and 4 sub-apertures.
3.2.3 Quantitative comparison
We calculate the quality metrics under different data conditions, with the results presented in Tables 2 to 5. Figures 14 visualizes these results using curve plots. The data reveal that all existing methods perform worse across all metrics compared to DL-based methods with cross-modal supervision, with particularly notable differences in the IoU and CE metrics. This disparity is largely due to modality differences between the output images and the ground truth data. And among the proposed DL-based methods, CMAR-Net consistently achieves the best performance across all metrics. The only exception occurs under the condition of 4 sub-apertures and an SNR of 5 dB, where CMAR-Net’s SSIM is slightly lower than the method using UNet3D as the backbone. We believe this is due to the UNet3D’s tendency to output higher pixel values (closer to 1). However, this behavior in fact leads to a loss of structural information, exhibiting poor generalization of the network, a conclusion that will be confirmed in real-world experiments.
In summary, compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 91.23% in CE, 80.28% in PSNR and 39.32% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 22.54% in IoU, 52.84% in CE, 24.32% in PSNR and 13.09% in SSIM. These results demonstrate that CMAR-Net can reconstruct more realistic images than other methods.
Table 2: IoU $\uparrow$ comparison across all methods on simulated data.
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP[35] 0.051 0.055 0.045 0.077 0.067 0.051 0.064 0.047 0.055
CS[34] 0.032 0.036 0.031 0.059 0.025 0.031 0.046 0.027 0.037
SACNet[23] 0.078 0.081 0.069 0.079 0.075 0.073 0.075 0.078 0.074
UNet3D[22] 0.031 0.030 0.041 0.027 0.041 0.030 0.038 0.034 0.046
SACNet+CM 0.653 0.744 0.269 0.717 0.755 0.434 0.558 0.462 0.592
UNet3D+CM 0.639 0.638 0.601 0.639 0.604 0.553 0.593 0.586 0.545
CMAR-Net(Ours) 0.750 0.744 0.713 0.745 0.726 0.689 0.694 0.670 0.626
Table 3: CE $\downarrow$ comparison across all methods on simulated data.
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP 1.624 1.685 1.637 1.625 1.640 1.636 1.656 1.629 1.654
CS 1.634 1.647 1.637 1.622 1.631 1.632 1.643 1.640 1.638
SACNet 1.307 1.294 1.091 1.357 1.196 1.263 1.283 1.332 0.992
UNet3D 0.491 0.503 0.593 0.511 0.545 0.565 0.526 0.533 0.607
SACNet+CM 0.205 0.171 0.530 0.215 0.153 0.371 0.464 0.498 0.387
UNet3D+CM 0.167 0.174 0.187 0.154 0.195 0.171 0.154 0.179 0.186
CMAR-Net(Ours) 0.099 0.108 0.096 0.112 0.109 0.104 0.119 0.111 0.139
[…]
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP 8.555 8.393 8.523 8.557 8.518 8.532 8.473 8.553 8.475
CS 8.538 8.502 8.530 8.573 8.548 8.549 8.511 8.525 8.527
SACNet 8.798 8.796 8.799 8.734 8.801 8.723 8.774 8.744 8.815
UNet3D 9.220 9.193 9.044 9.164 9.150 9.028 9.176 9.150 9.034
SACNet+CM 13.213 14.540 10.477 13.739 14.711 11.672 10.470 10.648 11.417
UNet3D+CM 13.183 13.143 12.964 13.409 12.806 13.157 13.333 12.971 12.897
CMAR-Net(Ours) 16.446 16.212 16.236 16.107 15.822 15.774 15.384 15.448 14.329
[…]
Methods
Number of
sub-aperture
12 8 4
SNR 30dB 15dB 5dB 30dB 15dB 5dB 30dB 15dB 5dB
BP 0.612 0.574 0.612 0.580 0.584 0.604 0.574 0.605 0.584
CS 0.626 0.603 0.628 0.591 0.637 0.628 0.589 0.626 0.606
SACNet 0.626 0.613 0.595 0.615 0.608 0.588 0.585 0.574 0.486
UNet3D 0.447 0.452 0.484 0.460 0.477 0.476 0.480 0.457 0.503
SACNet+CM 0.667 0.718 0.695 0.696 0.724 0.698 0.486 0.536 0.559
UNet3D+CM 0.763 0.762 0.750 0.774 0.744 0.757 0.766 0.753 0.746
CMAR-Net(Ours) 0.820 0.811 0.796 0.805 0.796 0.776 0.782 0.778 0.737
Figure 14: Image quality metrics-SNR curves for different methods on simulated data.
3.2.4 Feature space interpolation
Figure 15 shows the results of the feature space interpolation experiment. In this experiment, we perform linear interpolation between feature vectors extracted by the encoder structure in CMAR-Net, and observe how these interpolated features affect the generated output. The green boxes in the figure represent the output vehicle images of the training type. The images between each pair of green boxes represent the output results of the interpolated feature space. As shown in the figure, linear interpolation between two feature vectors leads to smooth and meaningful nonlinear transitions in the image space[36, 37]. Notably, the variations in body height and tail features of the interpolated images do not align with any specific vehicle type from the training data. This indicates that the network preserves distinct vehicle features during interpolation, rather than averaging them into an ”average” representation[38]. We believe this is a superior property of the reconstruction network, as it highlights the model’s ability to generate diverse outputs. This shows the model’s flexibility and capability to produce realistic and varied vehicle reconstructions.
Figure 15: Feature space interpolation results.
3.3 Results for real-world data
3.3.1 Real-world dataset
The real-world data used in this study is sourced from the GOTCHA Circular SAR dataset, collected and released by the United States Air Force Laboratory[39]. This dataset captures a scene with various civilian vehicles using a radar with the center frequency of 9.6 GHz and the bandwidth of 640 MHz. The data is acquired in circular SAR mode through 8 passes at different altitudes, with elevation angles of [45.66°, 44.01°, 43.92°, 44.18°, 44.14°, 43.53°, 43.01°, 43.06°]. The combination of circular observation aperture and altitude diversity enabled 3D imaging of the entire scene. We select data from two SUVs and two sedans. Figure 16 illustrates the scene, radar passes, and vehicle photos.
Figure 16: Real-world data description. (a) Data collection scene. (b) Radar flight passes. (c) Photos of test vehicles[39].
3.3.2 Imaging results
Figure 17 to 21 present the imaging results on real-world data with varying number of sub-apertures. The BP method produces results heavily affected by noise and side lobes, which can hardly be recognized. The CS method alleviates these issues, but the reconstructions still appear as chaotic point clouds, leaving vehicle features unrecognizable, especially when using highly sparse data. For example, when the number of sub-apertures is 4, traditional methods completely fail, producing totally disordered point clusters without any discernible or meaningful vehicle features.
In contrast, radar image-supervised DL methods can reconstruct vaguely discernible vehicle contours. However, as the number of apertures decreases, these contours deteriorates. For example, when the number of sub-apertures is reduced to 4, SACNet and UNet3D show significant omissions in reconstructing the roof of Jeep. After integrating cross-modal supervision, SACNet and UNet3D exhibit weak generalization when processing real-world data, struggling to accurately reconstruct the vehicle’s geometry. Particularly, when the number of sub-apertures is 4, SACNet fails completely and cannot output any visible features.
Figure 17: Imaging results of different methods using real-world data. Imaging using 12 sub-apertures
Figure 18: Imaging results of different methods using real-world data. Imaging using 10 sub-apertures
Figure 19: Imaging results of different methods using real-world data. Imaging using 8 sub-apertures
Figure 20: Imaging results of different methods using real-world data. Imaging using 6 sub-apertures
Figure 21: Imaging results of different methods using real-world data. Imaging using 4 sub-apertures
[…]
Methods Number of sub-aperture
12 10 8 6 4
BP 0.62/8.64 0.59/8.65 0.54/8.39 0.53/8.37 0.52/8.37
CS 0.50/8.68 0.49/8.69 0.45/8.63 0.43/8.61 0.40/8.60
SACNet 0.40/9.07 0.43/9.25 0.50/9.05 0.57/8.91 0.61/8.91
UNet3D 0.53/9.04 0.55/9.11 0.55/9.11 0.52/9.25 0.50/9.40
SACNet+CM 0.70/8.87 0.70/8.84 0.67/10.73 0.66/12.11 0.65/11.87
UNet3D+CM 0.68/11.65 0.69/11.98 0.68/11.97 0.69/12.03 0.71/12.29
w/o PRP unit 0.71/13.27 0.70/12.88 0.69/12.72 0.70/12.61 0.70/12.63
w/o data aug 0.72/13.31 0.70/12.47 0.69/13.04 0.68/13.08 0.68/12.90
CMAR-Net(Ours) 0.76/15.19 0.76/15.31 0.76/15.59 0.75/15.78 0.75/15.81
[…]
Methods Number of sub-aperture
12 10 8 6 4
BP 0.07/1.58 0.08/1.58 0.11/1.67 0.13/1.68 0.14/1.68
CS 0.03/1.57 0.04/1.57 0.05/1.59 0.05/1.60 0.07/1.60
SACNet 0.02/0.62 0.03/0.56 0.02/0.70 0.02/0.91 0.02/1.10
UNet3D 0.01/0.62 0.01/0.61 0.01/0.58 0.02/0.52 0.02/0.47
SACNet+CM 0.18/0.81 0.19/0.84 0.20/0.41 0.20/0.27 0.20/0.30
UNet3D+CM 0.53/0.32 0.57/0.29 0.56/0.28 0.57/0.28 0.60/0.25
w/o PRP unit 0.57/0.20 0.57/0.22 0.56/0.20 0.57/0.23 0.57/0.22
w/o data aug 0.55/0.23 0.55/0.29 0.48/0.22 0.45/0.20 0.45/0.21
CMAR-Net(Ours) 0.72/0.14 0.72/0.13 0.71/0.12 0.69/0.11 0.68/0.11
Figure 22: Image quality metrics-number of sub-aperture curves for different methods on real-world data.
Tables 6-7 and Figure 22 show the evaluation metrics. Compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 89.36% in CE, 75.83% in PSNR and 47.85% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 86.42% in IoU, 67.99% in CE, 38.81% in PSNR and 10.75% in SSIM.
Moreover, compared to simulated data, other DL methods experience a significant decline in performance, with an average drop of 12.17% across all metrics. In contrast, CMAR-Net maintains a high score, with only a 4.66% drop, significantly outperforming other methods and demonstrating its superior generalization ability. This is attributed to the PRP unit and data augmentation strategy, as further validated by the ablation experiments.
3.3.3 Ablation study
The 7th and 8th sets of data in the Figure 17-21 and table 6-7 show the imaging results and evaluation metrics after ablating the PRP unit and data augmentation strategy. Without the PRP unit, the vehicle images become blurry and suffer greater pixel loss as the number of apertures decreases. And in the absence of data augmentation, the network consistently reconstructs all vehicle type as sedan type, revealing severe overfitting. The PRP unit design improves SSIM, PSNR, IoU, and CE by 8.25%, 21.25%, 24.15%, and 42.82%, respectively. Data augmentation contributes improvements of 8.65%, 19.92%, 42.94%, and 46.75%. These results clearly demonstrate the effectiveness of both the PRP unit and data augmentation strategy.
4 Conclusion
In this paper, we propose a DL-based multi-baseline SAR 3D reconstruction method that utilizes only sparse aperture data. We introduce the concept of cross-modal learning and design a deep neural network tailored for SAR 3D reconstruction—CMAR-Net. By leveraging differentiable rendering, CMAR-Net can be effectively trained, addressing the key challenge of supervising 3D imaging network with 2D optical images. Extensive experiments on both simulated and real-world datasets show that CMAR-Net significantly outperforms all existing baseline methods, including traditional imaging algorithms and DL-based approaches, in terms of image quality and generalization. Moreover, the existing DL-based SAR reconstruction methods, adopting our proposed cross-modal learning paradigm, experience a significant performance improvement, which highlights the superiority and versatility. Last but not least, our method eliminates the need for time-consuming preprocessing of full-aperture data. Instead, it only needs 2D images obtained through computer rendering or camera capture. This significantly reduces the difficulty of dataset construction, greatly enhancing the practicality of SAR 3D reconstruction technology.
Our approach offers new insights for 3D SAR researchers and is expected to comprehensively improve the research landscape of SAR 3D imaging. The technique has the potential to be applied to fields such as aerial remote sensing, terrain mapping, and obstacle detection. In the future, researchers can expand the scope to 3D reconstruction of other scenes (e.g., buildings, streets), or explore fusion with other heterogeneous data (e.g., infrared data), and so on.